# Iterative Partition Search Variational Quantum Algorithm for Solving Shortest Vector Problem

### v1

@@ Abstract
The Partition Search Algorithm (PSA) and the Iterative Quantum Optimization with an Adaptive Problem (IQOAP) framework are two existing Variational Quantum Algorithms (VQAs) for solving the Shortest Vector Problem (SVP), but both suffer from certain limitations. In this work, we proposed the Iterative Partition Search Algorithm (IPSA), which is a targeted synthesis and refinement of these preceding methods. Our algorithm inherits the core idea of “partitioning to circumvent the zero vector” from PSA and the “iterative lattice basis reduction” framework from IQOAP. A key feature of IPSA is the “1-tailed search spaces”, which can be viewed as a highly constrained variant of PSA’s partitioning strategy, specifically designed for optimal performance within IQOAP’s iterative structure. We supplant IQOAP’s fixed iteration count with a dynamic, stack-managed process and substitute a more expressive and shallower circuit structure for its original ansatz. Crucially, the 1-tailed design fundamentally ensures that every successful VQA execution yields an effective lattice basis update, thereby eliminating the issue of ineffective iterations in IQOAP. This evolutionary path of refinement allows IPSA to overcome the drawbacks of its predecessors precisely. Numerical simulations on 4- to 6-dimensional SVP instances demonstrate that IPSA achieves at least a 73% improvement in success rate for finding optimal solutions and over a 35% improvement in average solution quality compared with the methods above while maintaining comparable total circuit depth.
@@ I  Introduction
The Shortest Vector Problem (SVP) is an NP-hard problem in lattice theory, underpinning the security of numerous post-quantum cryptographic schemes. Lattice is defined as a discrete additive subgroup of ${\mathbb{R}}^{n}$, typically generated by a basis matrix. These basis vectors form a lattice through their integer linear combinations. SVP seeks to identify the shortest non-zero vector within a lattice structure. Classical algorithms such as enumeration [1, 2, 3] and sieving [4, 5, 6] can solve SVP but require substantial computational costs. Quantum algorithms offer promising alternatives by exploiting Grover’s search [7] and quantum walks [8, 9, 10, 11] to achieve significant speedups. Recent quantum advances have reduced sieving complexity to $2^{0.2563n+o(n)}$ [10] and demonstrated quadratic speedups for enumeration method [11].
@@
However, these quantum advantages typically require fault-tolerant quantum computers, which remain beyond current technological capabilities. Present-day Noisy Intermediate-Scale Quantum (NISQ) devices are characterized by limited qubit counts and inherent noise, motivating the development of Variational Quantum Algorithms (VQAs) [12, 13, 14, 15, 16]. VQA represents hybrid quantum-classical computing paradigms that employ Parameterized Quantum Circuits (PQC), described by parameters $\bm{\theta}$, to prepare trial states $|\psi(\bm{\theta})\rangle$. A classical optimizer iteratively adjusts $\bm{\theta}$ to minimize a cost function $C(\bm{\theta})=\langle\psi(\bm{\theta})|H|\psi(\bm{\theta})\rangle$, where $H$ represents a Hamiltonian encoding the optimization problem. With a shallower circuit and inherent noise tolerance [17], VQA is well-suited for NISQ devices, with prominent examples ranging from foundational algorithms like the Variational Quantum Eigensolver (VQE) [18] and the Quantum Approximate Optimization Algorithm (QAOA) [19], to broader applications in areas like quantum federated learning [20] and quantum neural networks [21, 22, 23].
@@
Investigating VQA-based solvers for SVP is a meaningful endeavor, as it helps assess the potential threat that NISQ devices pose to post-quantum cryptography. Applying VQA to SVP involves three key aspects of consideration, and several existing studies have begun to address each of these points.
@@
First, mapping the infinite coefficients search space $\mathbb{Z}^{n}$ of SVP onto finite qubit count requires binary encoding strategy [24] and the establishment of bounded search ranges [25, 26], where the range boundaries are typically guided by classical preprocessing algorithms such as Lenstra–Lenstra–Lovász (LLL) [27] and Hermite-Korkin-Zolotarev (HKZ) [1]. Recently, Zhu, Joseph et al. proposed the Iterative Quantum Optimization with an Adaptive Problem (IQOAP) framework for solving SVP [28], achieving qubit requirement reduction compared with conventional boundary constraint methods [25, 26].
@@
Second, the straightforward Hamiltonian construction for the vector norm objective leads to a trivial all-zero ground state, necessitating sophisticated approaches to avoid this trivial solution. Standard avoidance strategies, such as introducing penalty terms and redesigning the cost function, come with certain costs due to the difficulty in tuning penalty coefficients and the requirement for additional qubits [25, 29, 30]. The Partition Search Algorithm (PSA) [31], originally proposed and validated in adiabatic quantum computing, offers an elegant approach to bypass this trivial solution issue without incurring any additional costs. Although initially developed within the adiabatic quantum computing framework, PSA demonstrates equivalent efficacy in VQA for avoiding trivial solutions.
Third, the design of PQC plays a crucial role in algorithm effectiveness. Standard ansatz like QAOA, while effective for certain optimization problems [32, 33], typically require deep circuit architectures for SVP due to the all-to-all connectivity inherent in the problem Hamiltonian [28, 30, 26]. This limitation motivates the adoption of alternative circuit structures, such as the Hardware-Efficient Ansatz (HEA) [34, 35, 36, 37, 38], which can provide strong expressive power with shallower circuits. Recently, we have also noted other mitigation strategies, such as using Fixed-Angle QAOA variants [39, 40, 41] or leveraging classical post-processing [42].
However, both PSA and IQOAP, despite their contributions, exhibit significant room for improvement. The complex optimization landscape of PSA often leads to convergence on low-quality suboptimal solutions, while its qubit requirement can be further reduced. On the other hand, IQOAP does not inherently avoid the trivial all-zero solution, and its effectiveness is constrained by a weak circuit, a fixed number of iterations, and the potential for ineffective iterations that waste computational resources.
To address these issues, we propose the Iterative Partition Search Algorithm (IPSA), which is designed as a targeted synthesis and refinement of these two preceding methods. IPSA inherits the core idea of “partitioning to circumvent the zero vector” from PSA and the “iterative lattice basis reduction” framework from IQOAP. A key feature of IPSA is the “1-tailed search space”, which acts as a highly constrained variant of PSA’s strategy, specifically engineered for optimal performance within this iterative structure. We replace the fixed iteration count with a dynamic, stack-managed process and substitute the initial ansatz with a more expressive circuit structure. Crucially, this 1-tailed design ensures that every successful VQA execution yields an effective lattice basis update, thereby fundamentally solving the issue of ineffective iterations. These refinements enable IPSA to achieve superior success rates and solution quality without a significant increase in computational cost overhead, thus advancing the application of VQA for solving SVP.
@@
The remainder of this paper is organized as follows. In Section II, we review two existing methods, the PSA and the IQOAP. In Section III, we present our proposed IPSA and detail its key components, including the 1-tailed search spaces, the stack-based management of search partitions, and the improved quantum circuit design. Section IV reports the results of our numerical simulations, where we benchmark the three algorithms on two distinct sets of SVP instances. In Section V, we provide our conclusions and a brief outlook for future applications.
@@ II  Foundation: Partition Search and Iterative Optimization Framework
The PSA [31] addresses the zero-vector issue without incurring additional computational overhead. It partitions the coefficient space of basis vector combinations into non-overlapping regions and solves the SVP independently within each region. For an $n$-dimensional lattice, the coefficient space is divided into regions $X_{1},\ldots,X_{n}$, where each region $X_{i}$ is defined as:
@@
The ground state of each $H_{i}$ yields the shortest vector within the corresponding partition, and the global minimum among all partitions provides the SVP solution. The total qubit requirement $M$ depends on the sum of bits needed to represent each coefficient $m$, which is determined by the preprocessing of the lattice basis. The analysis in Ref. [25] shows that for an HKZ-reduced basis, the requirement is $M=\mathcal{O}(n\log n)$, while Ref. [26] concludes that using an LLL-reduced basis requires $M=n(n+1)$ qubits. Although PSA effectively circumvents the trivial solution, its complex optimization landscape poses a significant risk of convergence on low-quality suboptimal solutions. Moreover, its implementation via boundary constraint methods still requires a substantial number of qubits, presenting opportunities for further qubit count reduction.
To reduce the qubit utilization in boundary constraint methods, the IQOAP framework [28] performs VQA optimization within a smaller, fixed coefficient range, leveraging the fact that even suboptimal solutions can progressively reduce the lattice basis through iterative refinement. For example, in 4-dimensional SVP instances, each $x_{r}$ is encoded using only 22 qubits via an encoding like:
@@
The IQOAP framework improves the basis quality through an iterative refinement process, which is outlined in Algorithm 1. In each iteration, if the VQA finds a vector $\bm{v}$ shorter than a current basis vector $\bm{b}_{j}$, then $\bm{b}_{j}$ is replaced by $\bm{v}$ (provided the lattice remains unchanged). Following the original study, this process is repeated for a fixed number of times. For instance, the authors state that 50 iterations suffice to solve 4-dimensional SVP instances with a high probability [28].
@@
3:Solve SVP with QAOA ansatz, obtaining the result $\bm{v}$.
[…]
5:the lattice remains unchanged.
6:If there are multiple vectors that can be replaced in
7:Step 3, choose the longest one.
[…]
While PSA can potentially benefit from IQOAP’s qubit reduction approach, IQOAP itself suffers from several critical limitations. First, it employs a fixed iteration count and single-layer QAOA ansatz with constrained parameters, which severely restricts the quantum circuit’s expressiveness for higher-dimensional SVP instances. Second, if the vector $\bm{v}$ obtained in Step 2 cannot replace any basis vector $\bm{b}_{j}$ without altering the lattice structure, the entire iteration becomes a wasted computation.
To address these limitations while preserving the advantages of both approaches, we propose the IPSA, which combines PSA’s partition strategy with an advanced iterative framework. We employ a stack data structure to iterate across different partitions dynamically, addressing IQOAP’s fixed iteration count limitation. Additionally, we replace IQOAP’s QAOA ansatz with the HEA, which offers greater expressiveness with reduced circuit depth. Furthermore, our proposed 1-tailed search spaces technique not only achieves greater qubit reduction and improved solution quality compared with IQOAP but also eliminates wasted computations that may occur in IQOAP.
@@ III  Iterative Partition Search Algorithm
In this section, we present the overall framework of the IPSA, including the specific design of the parameterized quantum circuit and the detailed analysis of 1-tailed search spaces.
IPSA operates on an iterative principle, using a stack to manage the search partitions dynamically. This stack-based strategy is designed to prioritize re-solving SVP within smaller partitions immediately after a basis vector is updated. The goal of this strategy is to ensure the basis is well-reduced (i.e., composed of shorter, more orthogonal vectors) before proceeding to larger partitions, thereby aiming to reduce the overall computational cost. The complete algorithm is outlined in Algorithm 2.

### v2

@@ Abstract
The Partition Search Algorithm (PSA) and Iterative Quantum Optimization with an Adaptive Problem (IQOAP) are leading variational quantum algorithms for solving Shortest Vector Problem (SVP). However, each has limitations that restrict its practical impact. IQOAP suffers from ineffective iterations that fail to update the lattice basis, whereas PSA’s static partitioning leads to oversized search spaces. In this work, we propose the Iterative Partition Search Algorithm (IPSA), which systematically addresses these drawbacks by integrating a “1-tailed search spaces” with a dynamic, stack-managed iterative process. Specifically, the “1-tailed” strategy ensures that every successful execution yields an effective lattice basis update, thereby eliminating the ineffective iterations associated with IQOAP. Concurrently, the dynamic iterative process reduces the required qubit count, thereby avoiding the limitation of an oversized search space inherent to PSA. We validate IPSA on the Baihua superconducting quantum processor via the Quafu platform. Small-scale real hardware experiments demonstrate that, compared to PSA, IPSA achieves a 14-fold increase in success rate at a cost of less than double the total circuit depth. Conversely, compared to IQOAP, IPSA reduces the total circuit depth by 82.7% while achieving approximately 2.5 times its success rate. Furthermore, we also conduct numerical simulations whose results are in good agreement with the experimental findings and extend our analysis.
@@ I  Introduction
The Shortest Vector Problem (SVP) is an NP-hard problem in lattice theory, underpinning the security of numerous post-quantum cryptographic schemes. Lattice is defined as a discrete additive subgroup of ${\mathbb{R}}^{n}$, typically generated by a basis matrix. These basis vectors form a lattice through their integer linear combinations. SVP seeks to identify the shortest non-zero vector within a lattice structure. Classical algorithms such as enumeration [kannan1983improved, fincke1985improved, gama2010lattice] and sieving [ajtai2001sieve, micciancio2010faster, becker2016new] can solve SVP but require substantial computational costs. Quantum algorithms offer promising alternatives by exploiting Grover’s search [grover1996fast] and quantum walks [laarhoven2015finding, chailloux2021lattice, bonnetain2023finding, aono2018quantum] to achieve significant speedups. Recent quantum advances have reduced sieving complexity to $2^{0.2563n+o(n)}$ [bonnetain2023finding] and demonstrated quadratic speedups for enumeration method [aono2018quantum].
@@
However, these quantum advantages typically require fault-tolerant quantum computers, which remain beyond current technological capabilities. Present-day Noisy Intermediate-Scale Quantum (NISQ) devices are characterized by limited qubit counts and inherent noise, motivating the development of Variational Quantum Algorithms (VQAs) [cerezo2021variational, lubasch2020variational, liu2021variational, wang2021variational, bravo2023variational]. VQA represents hybrid quantum-classical computing paradigms that employ parameterized quantum circuits, described by parameters $\bm{\theta}$, to prepare trial states $|\psi(\bm{\theta})\rangle$. A classical optimizer iteratively adjusts $\bm{\theta}$ to minimize a cost function $C(\bm{\theta})=\langle\psi(\bm{\theta})|H|\psi(\bm{\theta})\rangle$, where $H$ represents a Hamiltonian encoding the optimization problem. With a shallower circuit and inherent noise tolerance [mcclean2016theory], VQA is well-suited for NISQ devices, with prominent examples ranging from foundational algorithms like the Variational Quantum Eigensolver (VQE) [peruzzo2014variational] and the Quantum Approximate Optimization Algorithm (QAOA) [farhi2014quantum], to broader applications in areas like quantum federated learning [song2024quantum] and quantum neural networks [wu2024quantum, li2025efficient, su2025topology].
@@
Investigating VQA-based solvers for SVP is a meaningful endeavor, as it helps assess the potential threat that NISQ devices pose to post-quantum cryptography. Applying VQA to SVP involves two key aspects of consideration, and several existing studies have begun to address each of these points.
@@
First, mapping the infinite coefficients search space $\mathbb{Z}^{n}$ of SVP onto finite qubit count requires binary encoding strategy [joseph2021two] and the establishment of bounded search ranges [albrecht2023variational, zheng2025quantum], where the range boundaries are typically guided by classical preprocessing algorithms such as Lenstra–Lenstra–Lovász (LLL) [lenstra1982factoring] and Hermite-Korkin-Zolotarev (HKZ) [kannan1983improved]. Recently, Zhu, Joseph et al. proposed the Iterative Quantum Optimization with an Adaptive Problem (IQOAP) framework for solving the SVP [zhu2022iterative], achieving qubit requirement reduction compared with conventional boundary constraint methods [albrecht2023variational, zheng2025quantum].
@@
Second, the straightforward Hamiltonian construction for the vector norm objective leads to a trivial all-zero ground state, necessitating sophisticated approaches to avoid this trivial solution. Standard avoidance strategies, such as introducing penalty terms and redesigning the cost function, come with certain costs due to the difficulty in tuning penalty coefficients and the requirement for additional qubits [albrecht2023variational, mizuno2024quantum, barbera2024finding]. The Partition Search Algorithm (PSA) [yamaguchi2022annealing], originally proposed and validated in adiabatic quantum computing, offers an elegant approach to bypass this trivial solution issue without incurring any additional costs. Although initially developed within the adiabatic quantum computing framework, PSA demonstrates equivalent efficacy in VQA for avoiding trivial solutions.
However, despite their respective contributions, both PSA and IQOAP possess inherent limitations that hinder their practical performance. For PSA, its static partitioning strategy requires a high number of qubits, which in turn generates an oversized search space. This vast search space frequently traps the VQA in suboptimal local minima due to the complex optimization landscape. IQOAP, on the other hand, does not incorporate any mechanism to circumvent the trivial all-zero solution. Furthermore, its iterative structure is prone to ineffective iterations, in which a successfully found shorter vector is discarded because it fails to meet the requirements for a lattice basis update, thereby wasting computational resources.
To address these issues, we propose the Iterative Partition Search Algorithm (IPSA), which systematically addresses these drawbacks by integrating “1-tailed search spaces” into a dynamic, stack-managed iterative process. The dynamic iterative process directly addresses the limitations of PSA’s static approach. By enabling the use of smaller, successive search spaces, it fundamentally avoids the vast and complex optimization landscape that plagues PSA. Concurrently, the “1-tailed search spaces” resolve the primary drawbacks of IQOAP. It inherently avoids the trivial all-zero solution and ensures that every successfully found shorter vector meets the requirements for a lattice basis update, thereby eliminating ineffective iterations. As a result, IPSA demonstrates a superior trade-off between solution quality and computational cost compared to its predecessors, further advancing the application of VQA for solving the SVP.
@@
The remainder of this paper is organized as follows. In Section II, we briefly review two existing methods, the PSA and the IQOAP. In Section III, we present our proposed IPSA and detail its key components, including the 1-tailed search spaces and the stack-managed iterative process of search partitions. Section IV reports the results of our experiments on real quantum hardware. In Section V, we provide our conclusions and a brief outlook for future applications. A discussion justifying the selection of our parameterized quantum circuit is presented in Appendix A. Detailed information on the quantum processor’s parameters and topology is provided in Appendix B. Further numerical simulations are presented in Appendix C.
@@ II  Foundation: Partition Search and Iterative Optimization Framework
The PSA [yamaguchi2022annealing] addresses the zero-vector issue without incurring additional computational overhead. It partitions the coefficient space of basis vector combinations into non-overlapping regions and solves the SVP independently within each region. For an $n$-dimensional lattice, the coefficient space is divided into regions $X_{1},\ldots,X_{n}$, where each region $X_{i}$ is defined as:
@@
The ground state of each $H_{i}$ yields the shortest vector within the corresponding partition, and the global minimum among all partitions provides the SVP solution. The total qubit requirement $M$ depends on the sum of bits needed to represent each coefficient $m$, which is determined by the preprocessing of the lattice basis. The analysis in Ref. [albrecht2023variational] shows that for an HKZ-reduced basis, the requirement is $M=\mathcal{O}(n\log n)$, while Ref. [zheng2025quantum] concludes that using an LLL-reduced basis requires $M=n(n+1)$ qubits.
To reduce the qubit utilization in boundary constraint methods, the IQOAP framework [zhu2022iterative] performs VQA optimization within a smaller, fixed coefficient range, leveraging the fact that even suboptimal solutions can progressively reduce the lattice basis through iterative refinement. For example, in 4-dimensional SVP instances, each $x_{r}$ is encoded using only $2$ qubits via an encoding like:
@@
The IQOAP framework improves the basis quality through an iterative refinement process, which is outlined in Algorithm 1. In each iteration, if the VQA finds a vector $\bm{v}$ shorter than a current basis vector $\bm{b}_{j}$, then $\bm{b}_{j}$ is replaced by $\bm{v}$, provided the lattice remains unchanged. Following the original study, this process is repeated for a fixed number M of times.
@@
2:Output: The shortest vector in the final basis $B$.
3: while $iter<\text{M}$ do
4:   Solve SVP with QAOA ansatz, obtaining the result $\bm{v}$.
@@
7:   If there are multiple vectors that can be replaced in 8:Step 3, choose the longest one.
9:   $iter=iter+1$.
@@ III  Iterative Partition Search Algorithm
In this section, we present the Iterative Partition Search Algorithm (IPSA). We will detail its two core components: the “1-tailed search spaces” and the dynamic, stack-managed iterative process, which are designed to systematically address the limitations of previous methods. In addition, we will discuss the parameterized quantum circuit selected for its implementation on superconducting quantum processor.
III.1 1-Tailed Search Spaces
$~Y_{i}$ $=$ $\{(y_{1},\cdots,y_{i-1},1,0,\cdots,0)^{T}\in\mathbb{Z}^{n}\}$ (7)
$~\bm{v}=\sum_{j=1}^{i-1}y_{j}\bm{b}_{j}+\bm{b}_{i}$ (8)
This design, which fixes the coefficient of the $i$-th basis vector to 1, offers several crucial advantages. It reduces qubit requirements by eliminating the encoding for one coefficient. It also simplifies the cost function landscape, mitigating interference from other near-optimal vectors of very similar length. Most critically, since the $i$-th coefficient is fixed to 1, any solution $\bm{v}$ found in $Y_{i}$ can directly replace $\bm{b}_{i}$ if $\|\bm{v}\|<\|\bm{b}_{i}\|$ while preserving the lattice structure, as formalized in Theorem 1. The “1-tailed” strategy provides a crucial guarantee that the original IQOAP lacks: while IQOAP may successfully find a shorter vector, it cannot ensure the vector is valid for a basis update. This constrained partitioning, however, guarantees that any shorter solution found by the VQA is always valid for replacement, thereby fundamentally eliminating the problem of ineffective iterations.
Theorem 1
III.2 Stack-Managed Iterative Process
IPSA operates on an iterative principle, using a stack to manage the “1-tailed” search partitions dynamically. This stack-managed strategy is designed to prioritize re-solving the SVP within smaller partitions immediately after a basis vector is updated. This strategy try to ensure the basis is well-reduced (i.e., composed of shorter, more orthogonal vectors) before proceeding to larger partitions, thereby aiming to reduce the overall computational cost. The complete algorithm is outlined in Algorithm 2.
This dynamic, stack-managed strategy stands in contrast to PSA’s static approach. As previously discussed, the fixed PSA’s optimization landscape means that any convergence to a low-quality, suboptimal local minimum solution is worthless. However, IPSA’s dynamic stack transforms these solutions into meaningful steps. As shown in Algorithm 2 (Lines 6-8), a successful basis update causes the algorithm to re-search for the smallest partition with the updated basis. This part of IPSA differs from PSA, which performs only a single, fixed optimization.